# SoK: Runtime Integrity

### v2

@@ 4.2.2 Control Flow Bending (CFB)
| Device Type/Target | Scope | Overheads | ||||||||||||||
| Year | Scheme |
Embedded (bare-metal) |
Embedded (OS) |
High-end (User-space) |
High-end (Kernel) |
Mechanism | Sensitivity | System Support |
Data-Only |
ROP |
JOP |
Evidence Expressiveness |
Runtime |
Code Size |
Custom Hardware |
Network |
| Control Flow Integrity (CFI) Approaches | ||||||||||||||||
| 2013 | bin-CFI [86] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU | ✗ | ○ | ○ | - | ⚫ | ⚫ | ✗ | - |
| 2013 | CCFIR [87] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | ✗ | OS/MMU | ✗ | ○ | ○ | - | ⚫ | ⚫ | ✗ | - |
| 2014 | LLVM CFI [19] | ✗ | ✗ | ✓ | ✓ | SWI | ✗ | OS/MMU | ✗ | ✗ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2014 | KCoFI [88] | ✗ | ✗ | ✗ | ✓ | SWI | ✗ | MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2014 | MCFI [89] | ✗ | ✗ | ✓ | ✗ | SWI | CS | OS/MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2014 | RockJIT [137] | ✗ | ✗ | ✓ | ✗ | SWI | CS | OS/MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2015 | CCFI [90] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2015 | HAFIX [126] | ✓ | ✗ | ✗ | ✗ | ISA | Path | C-HW | ✗ | ● | ✗ | - | ⚫ | ⚫ | ⚫ | - |
| 2015 | CFCI [91] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU | ✗ | ○ | ○ | - | ⚫ | ⚫ | ✗ | - |
| 2015 | O-CFI [92] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | ✗ | OS/MMU+MPX | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2015 | CFI [82] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2015 | PathArmor [104] | ✗ | ✗ | ✓ | ✗ | SWI | Path | OS/MMU+LBR | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2015 | Lockdown [93] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | ✗ | OS/MMU | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2016 | TypeArmor [94] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU | ✗ | ✗ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2016 | FG-CFI [95] | ✗ | ✗ | ✗ | ✓ | SWI | ✗ | MMU | ✗ | ◐ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2016 | HCFI [123] | ✗ | ✗ | ✓ | ✗ | ISA | ✗ | OS+C-HW | ✗ | ● | ● | - | ⚫ | ⚫ | ⚫ | - |
| 2017 | PittyPAT [105] | ✗ | ✗ | ✓ | ✗ | SWI | Path | OS/MMU+PT | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2017 | GRIFFIN [124] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU+PT+TSX | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2017 | CFI-CaRE [133] | ✓ | ✗ | ✗ | ✗ | SWI+R/I | ✗ | TZ | ✗ | ● | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2017 | Intel CET [22] | ✗ | ✗ | ✓ | ✓ | ISA | ✗ | OS/MMU+CET | ✗ | ● | ○ | - | ⚫ | ⚫ | ✗ | - |
| 2018 | CFI [81] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU+PT | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2018 | SCFP [106] | ✓ | ✗ | ✗ | ✗ | SWI+C-HW | Path | C-HW | ✗ | ● | ◐ | - | ⚫ | ⚫ | ⚫ | - |
| 2018 | ARM BTI [96] | ✓ | ✓ | ✓ | ✓ | ISA | ✗ | BTI | ✗ | ✗ | ○ | - | ⚫ | ⚫ | ✗ | - |
| 2018 | PAC-RET [97] | ✓ | ✓ | ✓ | ✓ | ISA | ✗ | PA | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
| 2019 | OS-CFI [107] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | Path | OS/MMU+MPX+TSX | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
| 2019 | CFI-LB [108] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | CS | OS/MMU+TSX | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
| 2019 | PARTS [26] | ✗ | ✗ | ✓ | ✗ | SWI+ISA | ✗ | OS/MMU+PA | + | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2020 | RAI [109] | ✓ | ✗ | ✗ | ✗ | SWI+R/I | Path | MPU | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
| 2020 | Silhouette [127] | ✓ | ✗ | ✗ | ✗ | SWI+R/I | ✗ | MPU | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
| 2021 | VIP [110] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | Path | OS/MMU+MPK | + | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
| 2021 | PACStack [28] | ✗ | ✗ | ✓ | ✗ | SWI+ISA | ✗ | OS/MMU+PA | ✗ | ● | ✗ | - | ⚫ | ⚫ | ✗ | - |
| 2022 | TyPro [125] | ✗ | ✗ | ✓ | ✗ | SWI | ✗ | OS/MMU | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
| 2022 | PAL [27] | ✗ | ✗ | ✗ | ✓ | SWI+ISA | ✗ | PA+MMU | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2022 | PACTight [98] | ✗ | ✗ | ✓ | ✗ | SWI+ISA | ✗ | OS/MMU+PA | ✗ | ● | ● | - | ⚫ | ⚫ | ✗ | - |
| 2023 | FineIBT [23] | ✗ | ✗ | ✓ | ✓ | SWI+ISA | ✗ | CET+MMU | ✗ | ✗ | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2023 | SHERLOC [112] | ✓ | ✗ | ✗ | ✗ | SWI | Path | TZ+MTB+DWT | ✗ | ● | ◐ | - | ⚫ | ⚫ | ✗ | - |
| 2023 | TypeSqueezer [147] | ✗ | ✗ | ✓ | ✗ | SWI | Path | OS/MMU | ✗ | ✗ | ● | - | ⚫ | ⚫ | ✗ | - |
| 2024 | HEK-CFI [148] | ✗ | ✗ | ✗ | ✓ | ISA | ✗ | CET+MMU | ✗ | ● | ◐ | - | ⚫ | ⚫ | ✗ | - |
| Control Flow Attestation (CFA) Approaches | ||||||||||||||||
| 2016 | C-FLAT [34] | ✓ | ✗ | ✗ | ✗ | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2017 | LO-FAT [36] | ✓ | ✗ | ✗ | ✗ | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
| 2017 | ATRIUM [85] | ✓ | ✗ | ✗ | ✗ | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
| 2018 | LiteHAX [72] | ✓ | ✗ | ✗ | ✗ | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
| 2019 | DIAT [149] | ✓ | ✓ | ✗ | ✗ | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2019 | ScaRR [117] | ✗ | ✗ | ✓ | ✗ | SWI | -based | OS/MMU | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2019 | RIM [120] | ✓ | ✗ | ✗ | ✗ | C-HW | Path | C-HW | ● | ● | ✗ | ✗ | ? | |||
| 2020 | OAT [71] | ✓ | ✓ | ✗ | ✗ | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2020 | LAHEL [150] | ✓ | ✓ | ✗ | ✗ | C-HW | -based | C-HW | ◐ | ◐ | ⚫ | ⚫ | debug HW | |||
| 2020 | LAPE [151] | ✓ | ✗ | ✗ | ✗ | SWI+R/I | -based | MPU | ◐ | ◐ | ⚫ | ⚫ | ✗ | |||
| 2021 | Tiny-CFA [35] | ✓ | ✗ | ✗ | ✗ | SWI | -based | C-HW | ● | ● | ⚫ | ⚫ | ⚫ | |||
| 2021 | DIALED [73] | ✓ | ✗ | ✗ | ✗ | SWI | -based | C-HW | ● | ● | ⚫ | ⚫ | ⚫ | |||
| 2021 | ReCFA [121] | ✗ | ✗ | ✓ | ✗ | SWI+R/I | -based | OS+MPK | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2022 | GuaranTEE [122] | ✗ | ✗ | ✓ | ✗ | SWI | -based | Intel SGX | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2023 | ACFA [38] | ✓ | ✗ | ✗ | ✗ | C-HW | -based | C-HW | ● | ● | ✗ | ✗ | ⚫ | |||
| 2023 | ARI [37] | ✓ | ✗ | ✗ | ✗ | SWI | -based | TZ | ◐ | ◐ | ⚫ | ⚫ | ✗ | |||
| 2023 | BLAST [136] | ✓ | ✓ | ✗ | ✗ | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
| 2023 | ISC-FLAT [84] | ✓ | ✗ | ✗ | ✗ | SWI | -based | TZ | ● | ● | ⚫ | ⚫ | ✗ | |||
Legend:
✓ Has this feature,
✗ Lacks this feature,
- Feature is not applicable,
SWI: Software Instrumentation,
R/I: Randomization or Isolation,
ISA: Instruction Set Architecture,
MMU: Memory Management Unit,
C-HW: Custom Hardware,
OS: Operating System,
CS: Context Sensitive,
Path: Context- & path-sensitive
MPX: Intel Memory Protection eXtensions
LBR: Intel Last Branch Record
PT: Intel Processor Trace
TSX: Intel Transactional Synchronization Extension
TZ: ARM TrustZone,
CET: Intel Control-Flow Enforcement Tech.,
BTI: ARM Branch Target Identification,
PA: ARM Pointer Authentication,
MPU: Memory Protection Unit,
MPK: Intel Memory Protection keys,
MTB: ARM Macro Trace Buffer,
DWT: ARM Data Warchpoint and Trace,
⚫ Higher overhead,
? Feature required but cost not reported,
debug HW refers to reliance on prototyping/debug features not meant for device deployment.
@@ 4.2.3 Race Conditions
Additionally, WarpAttack [153] revealed that compiler optimizations can introduce double-fetch vulnerabilities, resulting in concurrency issues and TOCTOU, even with a strict static CFI policy that includes both forward and backward-edge protections. WarpAttack bypassed several CFI defenses, including LLVM-CFI [19], Lockdown [93], and MS-CFG [118]. To mitigate race conditions, contemporary CFI mechanisms rely on hardware support. For example, OS-CFI [107] and CFI-LB [108] utilize Intel TSX to safeguard intermediate values.
In CFA (and more broadly RA), resistance against TOCTOU attacks and race conditions often refers to achieving temporal consistency between when the executable binary is measured and when it is executed [154, 155, 156, 84, 38, 35]. Aside from modifications to code, the integrity of CFA evidence can be compromised by external interrupts that may stealthily modify the control flow path or the execution state, as shown and mitigated by ISC-FLAT [84].
@@ 5.1 Takeways
Considering question [Q1], posed in Section 1, our systematization presents several differences between CFI and CFA. The effectiveness of CFI mechanisms is intrinsically tied to the comprehensiveness and accuracy of a (statically-defined or dynamic) policy enforced locally. Most CFA techniques are policy-agnostic, passively monitoring execution to generate authenticated control flow reports. Contrary to CFI, CFA concerns convincing a remote party of trustworthy execution behavior, serving as a run-time analog to static attestation methods that prove the integrity of booted/loaded code. Thus, CFA reports are transmitted to a remote for analysis. These observations lead us to Takeaway 1.
@@
Regarding [Q2], we first examine CFA/CFI assumptions. Many CFI schemes assume the ability to apply WX on memory to preserve the software instrumentation (SWI) and avoid code injection. In Table I, this is apparent from user-space CFI schemes frequently relying on OS/MMU system support to enforce the WX policy. While CFA mechanisms need not impose WX, they must rely on an attestation RoT in to attest that reported runtime evidence is authentic (this includes code integrity and instrumentation, when applicable). Furthermore, unlike CFI, CFA requires network connectivity between and . Despite these differences, we also observe that state-of-the-art techniques for CFI and CFA intersect in their mechanisms for monitoring control flow events. For instance, many schemes utilize SWI as a mechanism while relying on hardware (whether commodity or custom) to protect or support SWI, as shown in Table I. Table I also shows that both CFI and CFA can eliminate some SWI using ISA-specific or custom hardware extensions. This is summarized in Takeaway 2.
@@
A common misconception/over-simplification that relates to [Q3] is that CFA’s entire purpose is to enable CFI checks to be outsourced to a resource-rich , avoiding CFI costs on . As extensively discussed in this systematization, CFA goals go beyond outsourcing CFI checks. As evidence of that, recent CFA methods have evolved to generate expressive (often lossless) control flow path evidence, as opposed to proving adherence to a locally enforced CFI policy. See Evidence Expressiveness column, in Table I. This is subsumed by Takeaway 3.
@@
Regarding [Q4], given their distinct security goals, the coexistence of CFI and CFA on the same platform could be possible if the performance overhead is acceptable in the target domain. We believe exploring hybrid approaches that combine the strengths of CFI and CFA to be an intriguing avenue for further research. A potential hybrid design might include CFI building blocks that can be elegantly incorporated into CFA reports. Considering that many state-of-the-art CFI offer fine-grained local ROP detection with low overheads (as seen in Scope and Overheads columns in Table I), a hybrid approach might implement CFI techniques for local ROP detection while utilizing CFA techniques for generating expressive evidence of path deviations due to JOP, DOP, and/or logic control bugs. However, CFI/CFA integration is not trivial, as differences in designs and system assumptions should be considered and can contribute to overheads. These observations are summarized in Takeaway 4.
@@ 5.2 Paths Foward
Demand for Stronger Threat Models. Currently considered threat models (in both CFI and CFA) can be limited in scope or may not adequately address the challenges posed by sophisticated adversaries (e.g., those capable of launching side-channel attacks). Next-generation mechanisms should consider stronger threat models to encompass new attack vectors that can lead to control-flow violations.
@@
CFA Support for Complex Software. The current landscape of CFA mechanisms primarily focuses on addressing the needs of simple, specialized, bare-metal embedded software (see column Device Type/Target in Table I). However, this limited scope poses challenges when it comes to applying these mechanisms to complex software scenarios with wider attack surfaces. To overcome this limitation, it is crucial to develop CFA mechanisms specifically tailored for complex software. Handling such a challenge not only helps align with zero-trust principles in demanding domains like cloud scenarios but also provides valuable insights into the effectiveness of CFA mechanisms compared to CFI mechanisms.
@@
CFA Evidence Verification & Practicality. The majority of CFA literature focuses on , assuming a able to interpret received evidence to detect attacks and identify root causes as long as the evidence is sufficiently expressive. Alas, there is a significant lack of concrete instances to substantiate postulated evidence analysis capabilities. Most of the CFA literature either leaves implementation as future work or implements simple remote checks based on received evidence, e.g., adherence to a CFG or emulated shadow stack (both of which could also be done locally by several CFI methods). Only two studies have explored alternative approaches. ZEKRA [165] suggests generating a zero-knowledge proof of CFG adherence for an untrusted , while RAGE [166] proposes training a Graph Neural Network (GNN) on previous runtime evidence for path verification. Yet, no prior work systematically demonstrates postulated benefits of CFA in terms of uncovering complex attacks (and root causes) based on remotely analyzed evidence.
Additionally, striking a balance between evidence expressiveness and overhead poses a challenge in achieving full-fledged CFA. Hashed paths compromise detailed runtime evidence in exchange for reduced storage and transmission costs. However, lossless path representations (and associated transmission to ) remain costly. As the complexity of the applications increases, the importance of expressiveness/cost trade-offs becomes more pronounced. Within this realm, promising avenues for future work include the development of mechanisms to reduce evidence storage and transmission costs while maintaining relevance and expressiveness.
CFI in Real-Time Systems and Other Niche. To the best of our knowledge, the majority of existing CFI mechanisms are not well-suited for real-time systems, highlighting the need for innovative approaches in this domain. Proposing effective CFI mechanisms for real-time systems entails addressing two critical challenges that are vital for ensuring both functional correctness and system safety.
The first challenge stems from the strict timing requirements inherent in real-time systems, particularly regarding intra-task timing. It is imperative to design CFI mechanisms that can operate within these constraints without compromising system performance. A potential avenue to tackle this challenge is the use of hardware-assisted branch monitors that reduce/eliminate code instrumentation. By parallelizing branch checks, hardware-based CFI could provide integrity without impacting intra-application delays.
@@
The second challenge involves rethinking the conventional approach of terminating an exploited application upon detecting a CFI violation, especially in domains such as autonomous systems. Abruptly terminating an application can introduce system instability or disruptions, posing risks to critical operations. Alternative strategies could be explored to recover from CFI violations to ensure system safety without unintended consequences. A promising direction to address this challenge is to design CFI schemes accommodating multi-variant execution that allows the containment of exploited applications while enabling the continuation of critical tasks.
We note that devising CFI that tackles both aforementioned challenges is non-trivial and requires more careful consideration. This includes accounting for factors such as portability, adaptability, and scalability.

### v3

@@ 4.2.3 Race Conditions
Additionally, WarpAttack [156] revealed that compiler optimizations could introduce double-fetch vulnerabilities, resulting in concurrency issues and TOCTOU, even with a strict static CFI policy that includes both forward and backward-edge protections. WarpAttack bypassed several CFI defenses, including LLVM-CFI [18], Lockdown [95], and MS-CFG [121]. To mitigate race conditions, contemporary CFI mechanisms rely on hardware support. For example, OS-CFI [109] and CFI-LB [110] utilize Intel TSX to safeguard intermediate values.
In CFA (and more broadly RA), resistance against TOCTOU attacks and race conditions often refers to achieving temporal consistency between when the executable binary is measured and when it is executed [157, 158, 159, 34]. Aside from modifications to code, the integrity of CFA evidence can be compromised by external interrupts that may stealthily modify the control flow path or the execution state, as shown and mitigated by ISC-FLAT [43].
@@ 5.1 Takeways
Considering question [Q1] posed in Section 1, this systematization presents several differences between CFI and CFA. The effectiveness of CFI mechanisms is intrinsically tied to the comprehensiveness and accuracy of a (statically-defined or dynamic) policy enforced locally. Most CFA techniques are policy-agnostic, passively monitoring execution to generate authenticated control flow reports. Contrary to CFI, CFA concerns convincing a remote party of trustworthy execution behavior, serving as a runtime analog to static attestation methods that prove the integrity of booted/loaded code. Thus, CFA reports are transmitted to a remote for analysis. These observations lead us to Takeaway 1.
@@
Regarding [Q2], we first examine CFA/CFI assumptions. Many CFI schemes assume the ability to apply WX on memory to preserve the integrity of IRMs and avoid code injection. In Table I, this is apparent from user-space CFI schemes frequently relying on OS/MMU system support to enforce the WX policy. While CFA mechanisms need not impose WX, they must rely on an attestation RoT in to attest that reported runtime evidence is authentic (this includes code integrity and instrumentation, when applicable). Furthermore, unlike CFI, CFA requires network connectivity between and . Despite these differences, we also observe that state-of-the-art techniques for CFI and CFA intersect in their mechanisms for monitoring control flow events. For instance, many schemes utilize IRMs via SWI as a mechanism while relying on hardware (whether commodity or custom) to protect or support their instrumentation, as shown in Table I. Table I also shows that both CFI and CFA can optimize SWI using specific ISA extensions. This is summarized in Takeaway 2.
@@
A common misconception/over-simplification that relates to [Q3] is that CFA’s entire purpose is to enable CFI checks to be outsourced to a resource-rich , avoiding CFI costs on . As extensively discussed in this systematization, CFA goals go beyond outsourcing CFI checks. As evidence of that, recent CFA methods have evolved to generate expressive (often lossless) control flow path evidence, as opposed to proving adherence to a locally enforced CFI policy (see Evidence Expressiveness column, in Table I). This is subsumed by Takeaway 3.
@@
Regarding [Q4], given their distinct security goals, the coexistence of CFI and CFA on the same platform could be possible if the performance overhead is acceptable in the target domain. We believe the exploration of approaches that combine the strengths of CFI and CFA to be an intriguing avenue for further research. A potential hybrid design might include CFI building blocks that can be elegantly incorporated into CFA reports. Considering that many state-of-the-art CFI offers fine-grained local ROP detection with low overheads (as seen in Scope and Overheads columns in Table I), a hybrid approach might implement CFI techniques for local ROP detection while utilizing CFA techniques for generating expressive evidence of path deviations due to JOP, and/or logic control bugs. Yet, CFI/CFA integration is non-trivial, as differences in designs and system assumptions should be considered and can contribute to overheads. As a first step in this direction, CFA+[120] recently proposed a mechanism that combines CFI and CFA to locally enforce specific targets for certain control flow transfers, relying on ARMv8.5-A landing pad instructions[98] while leveraging minimal instrumentation to record path information in reserved registers. These observations are summarized in Takeaway 4.
@@ 5.2 Paths Foward
Demand for Stronger Threat Models. Currently considered threat models (in both CFI and CFA) can be limited in scope or may not adequately address the challenges posed by sophisticated adversaries (e.g., those capable of launching side-channel attacks). Next-generation mechanisms could consider stronger threat models to encompass new attack vectors that can lead to control-flow violations.
@@
CFA Support for Complex Software. The current landscape of CFA mechanisms primarily focuses on addressing the needs of simple, specialized, bare-metal embedded software (see column Device Type/Target in Table I). This limited scope poses challenges when it comes to applying these mechanisms to complex software scenarios with wider attack surfaces. To overcome this limitation, it is crucial to develop CFA mechanisms specifically tailored for complex software.
@@
CFA Evidence Verification & Practicality. The majority of CFA literature focuses on , assuming a can interpret received evidence to detect attacks and identify root causes as long as the evidence is sufficiently expressive. Alas, there is a significant lack of concrete instances to substantiate postulated evidence analysis capabilities. Most of the CFA literature either leaves implementation as future work or implements simple remote checks based on received evidence, e.g., adherence to a CFG or emulated shadow stack (both of these could also be done locally by several CFI methods and still face the complex challenge of validating non-deterministic forward-edges). Only two studies have explored alternative approaches to simple remote checks. ZEKRA [168] suggests generating a zero-knowledge proof of CFG adherence for an untrusted , while RAGE [169] proposes training a Graph Neural Network (GNN) on previous runtime evidence for path verification. Yet, thus far, no prior work has concretely demonstrated CFA’s postulated benefits in uncovering complex attacks (and their root causes) based on remotely analyzed evidence. Additionally, striking a balance between evidence expressiveness and overhead poses a challenge in achieving full-fledged CFA. Hashed paths compromise detailed runtime evidence in exchange for reduced storage and transmission costs. However, lossless path representations (and associated transmission to ) remain costly. As the complexity of the applications increases, the importance of expressiveness/cost trade-offs becomes more pronounced. Within this realm, promising avenues for future work include the development of mechanisms to reduce evidence storage and transmission costs while maintaining relevance and expressiveness. In this direction, recent work in SpecCFA [148] proposes architectural support for -defined application-specific optimizations based on likely control flow sub-paths, enabling reduced storage/transmission costs without compromising evidence expressiveness.
CFI in Real-Time Systems and Other Niche. Most CFI proposals in Table I are not well-suited for real-time systems where strict timing requirements and execution integrity must be simultaneously maintained. Several recent CFI proposals focus on this gap [170, 171, 172, 173, 174]. InsectACIDE [170] uses architectural support (ARM TrustZone and MTB) to record control flow events without adding intra-task delays. During idle periods in between the execution of tasks, InsectACIDE uses the recorded information to perform security checks and locally detect control flow violations. FastCFI [171] also uses ARM features for tracing control flow transfers but relies on Field Programmable Gate Arrays (FPGA) to store and traverse a CFG according to these transfers. ECFI [174] proposes a mechanism for real-time Programmable Logic Controllers (PLCs) in which indirect branches are instrumented, and a PLC’s OS can schedule CFI checks. In ECFI, CFI checks are assigned a lower priority so that PLC tasks maintain their Worst-Case Execution Time (WCET). RECFISH [172] proposes CFI for ARM MCUs executing both bare-metal software or applications atop FreeRTOS. RECFISH utilizes SWI to insert trampolines to functionality enforcing indirect branch destinations according to function label sets or a shadow stack. When applied to FreeRTOS, RECFISH also saves the task’s state to the shadow stack to ensure that CFI-critical data cannot be overwritten during a context switch.
@@
Another challenge involves rethinking the conventional approach of terminating an exploited application upon detecting a CFI violation, especially in domains such as autonomous systems. Abruptly terminating an application can introduce system instability or disruptions, posing risks to critical operations. Current proposals for CFI in real-time systems generally focus on minimizing WCET while supporting local detection rather than delving into post-detection recovery strategies [171, 170, 172]. One alternative approach in ECFI [174] makes killing the violating process configurable via an optional flag and, by default, stores a log file of the violation details. Alternative strategies could be explored to recover from CFI violations and ensure system safety without unintended consequences. A promising direction to address this challenge is to design CFI schemes accommodating multi-variant execution that allows the containment of exploited applications while enabling the continuation of critical tasks. We note that devising CFI that tackles both aforementioned challenges is non-trivial and requires more careful consideration. This includes accounting for factors such as portability, adaptability, and scalability.
Acknowledgements
The authors sincerely thank the anonymous shepherd and the anonymous reviewers for their guidance and constructive criticism.